A boat is heading north as it crosses a wide river with a speed of 8.00 km/h relative to the water. The river has a uniform velocity of 6.00 km/h dues east. Determine the magnitude and direction of the boat's velocity with respect to an observer on the riverbank. (Note write all the possible direction of the motion)
1
Expert's answer
2020-11-30T14:51:48-0500

A boat must cross a river at direction "arcsin\frac 68\approx" 49° N of W.

Magnitude is "\sqrt{8^2-6^2}=\sqrt{28}\approx 5.29 ~\frac{km}{h}."